OpenStudy (anonymous):

Solve graphically x + y + 2 = 0 and 3x - 4y = 5. Write down the co-ordinates of the point of intersection of the two lines.

OpenStudy (anonymous):

3x - 4y = 5 4x + 4y = -8 7x = -3 x= -3/7 y = -11/7 co ordinates are ( -3/7, -11/7 )

OpenStudy (anonymous):

Thank, had a different answer I will fingd out what I did wrong. Thank you.